# The Born rule – 100 years ago and today

### v2

@@ 3.4.4 Domain of validity of the universal form
$\bullet$ As first observed by Heisenberg [32, p.25], the Born rule implies a tiny but positive probability that an electron bound to an atom is detected light years away from the atom: ”Das Resultat ist aber merkwürdiger, als es im ersten Augenblick den Anschein hat. Bekanntlich nimmt $\psi^{*}\psi$ exponentiell mit wachsendem Abstand vom Atomkern ab. Also besteht immer noch eine endliche Wahrscheinlichkeit dafür, das Elektron in sehr weitem Abstand vom Atomkern zu finden.” Therefore $|\psi(x)|^{2}$ is unlikely to be the exact probability density for being detected at $x$, as (BR-US) would require. This indicates that for observables with continuous spectrum, (BR-US) must be an idealization.
@@ 3.4.5 Domain of validity of the condensed form
The statistical interpretation based on the Born rule in the condensed form (BR-C) and in the POVM form (BR-POVM) are derivable from the DRP, and give a correct account of the actual experimental situation. They apply without idealization to results of arbitrary quantum measurements in the sense of the DRP. For a large number of examples see Neumaier & Westra [54, Section 6.2].
However, (BR-C) and (BR-POVM) still have the same limitations as the DRP, and cannot be applied to measurement of the kind discussed at the beginning of this subsection, such as macroscopic measurements or the measurements of spectral lines. For example, spectroscopy-based high precision measurements such as the gyromagnetic ratio of the electron are not quantum measurements but measurements of numerical parameters in specific quantum models; see the detailed discussion in Neumaier & Westra [54, Section 5.2].
@@ 3.5 What is missing in the foundations?
The above notion of what is must be distingushed from the theoretical ontology of certain interpretations of quantum mechanics: In the Copenhagen interpretation, what is is the classical world, delineated from the quantum world by a movable Heisenberg cut. In QBism (Fuchs [27, Section 3]), what is (in the present sense) is degraded to the world of a metaphysical society of agents governed by moral imperatives of how they should bet. In many world interpretations, a preferred ontological status is given to a postulated universal wave function, a theoretical figment of abstraction, whereas what is (in the present sense) is degraded to a particular world amidst infinitely many worlds, namely the single distinguished world that is our actual past, and from which we have to learn everything we can know about Nature, independent of any weight one might give to this world. In Bohmian mechanics, infinitely precise positions and the exact wave function are declared to have a preferred ontological status – a poor, theoretical substitute for the colorful what is that we experience. In each case, nothing real (in the sense of our what is) is in the quantum theory (of states and observables) itself, a state of affairs criticized by Einstein throughout his life.
@@
The work by Allahverdyan et al. [2] goes some way towards resolving the problem. But they have to equate at some points the formal probabilities of statistical mechanics (cf. Subsection 1.3) with measured probabilities, thus silently invoking the Born rule. This is done at the beginning of Section 11, where they say that ”the statistical interpretation of quantum mechanics emphasizes the idea that this theory, whether it deals with pure or mixed states, does not govern individual systems but statistical ensembles of systems”. But it describes statistical ensembles only if the formal probabilities are taken to be observed frequentist probabilities, which requires Born’s rule. Thus the derivation is circular. The problem is still present in their recent paper [3, p.254], where we read: ”we acknowledge that each individual run of the measurement produces a single macroscopic indication of the pointer of the apparatus which may be read or registered. This statement, although intuitive, is not a consequence of the mere quantum principles and has a macroscopic nature. We will accept its inclusion in quantum theory as a postulate (motivated by the macroscopic size of M)”. In fact, when they write that ”what we call ’quantum state’, whether pure or not, is in general an abstract mathematical object which allows us to make probabilistic predictions about the outcomes of any experiment”, they assume much more, namely the condensed form (BR-C) of Born’s rule! In effect they show that (BR-C) assumed for quantum statistical mechanics implies the Born rule for individual events.
@@
The main problem behind these difficulties is the need to find a rigorous derivation of a dissipative classical dynamics for a single pointer (or switch) without a silent recourse to the Born rule. A customary way to proceed (following Möhring & Smilansky [47]) is to start with the influence functional approach of Feynman & Vernon [24]. However, this approach has the Born rule built in into the derivation (p.122) of the basic functional integral representation of the transition probabilities.
A solution of the measurement problem is indeed impossible unless one has a precise mathematical definition of when and how a mathematical model of a single quantum detector produces a definite response.666 ”There is no consistent quantum measurement theory, unless an appropriate reinterpretation of what it means for an observable to have a definite value can be found.” (Busch & Lahti [14, p.375]) This is the reason why, 100 years after the inception of quantum mechanics in terms of states and observables, the quantum measurement problem is still unsolved. It is the missing piece in the foundations of quantum mechanics.

### v3

@@ 3.4.4 Domain of validity of the universal form
$\bullet$ As first observed by Heisenberg [36, p.25], the Born rule implies a tiny but positive probability that an electron bound to an atom is detected light years away from the atom: ”The result is more remarkable that it seems at first. For it is known that $\psi^{*}\psi$ decreases exponentially with increasing distance. Hence there is always a positive probability for finding the electron very far from the nucleus.”
(German original: ”Das Resultat ist aber merkwürdiger, als es im ersten Augenblick den Anschein hat. Bekanntlich nimmt $\psi^{*}\psi$ exponentiell mit wachsendem Abstand vom Atomkern ab. Also besteht immer noch eine endliche Wahrscheinlichkeit dafür, das Elektron in sehr weitem Abstand vom Atomkern zu finden.”) Therefore $|\psi(x)|^{2}$ is unlikely to be the exact probability density for being detected at $x$, as (BR-US) would require. This indicates that for observables with continuous spectrum, (BR-US) must be an idealization.
@@ 3.4.5 Domain of validity of the condensed form
The statistical interpretation based on the Born rule in the condensed form (BR-C) and in the POVM form (BR-POVM) are derivable from the DRP, and give a correct account of the actual experimental situation. They apply without idealization to results of arbitrary quantum measurements in the sense of the DRP. For a large number of examples see Neumaier & Westra [58, Section 6.2].
However, (BR-C) and (BR-POVM) still have the same limitations as the DRP, and cannot be applied to measurement of the kind discussed at the beginning of this subsection, such as macroscopic measurements or the measurements of spectral lines. For example, spectroscopy-based high precision measurements such as the gyromagnetic ratio of the electron are not quantum measurements but measurements of numerical parameters in specific quantum models; see the detailed discussion in Neumaier & Westra [58, Section 5.2].
@@ 3.5 What is missing in the foundations?
The above notion of what is must be distingushed from the theoretical ontology of certain interpretations of quantum mechanics: In the Copenhagen interpretation, what is is the classical world, delineated from the quantum world by a movable Heisenberg cut. In the consistent histories approach (Griffiths [34]), what is is the single history up to our present, which is just one of an uncountable collection of consistent histories about which theory only contributes uncheckable probability assignments to each of them. In many world interpretations, a preferred ontological status is given to a postulated universal wave function, a theoretical figment of abstraction, whereas what is (in the present sense) is degraded to a particular world amidst infinitely many worlds, namely the single distinguished world that is our actual past, and from which we have to learn everything we can know about Nature, independent of any weight one might give to this world. In Bohmian mechanics, infinitely precise positions and the exact wave function are declared to have a preferred ontological status – a poor, theoretical substitute for the colorful what is that we experience. In QBism (Fuchs [30, Section 3]), what is (in the present sense) is degraded to the world of a metaphysical society of agents governed by moral imperatives of how they should bet. In each case, nothing real (in the sense of our what is) is in the quantum theory (of states and observables) itself, a state of affairs criticized by Einstein throughout his life.
@@
The work by Allahverdyan et al. [3] goes some way towards resolving the problem. But they have to equate at some points the formal probabilities of statistical mechanics (cf. Subsection 1.3) with measured probabilities, thus silently invoking the Born rule. This is done at the beginning of Section 11, where they say that ”the statistical interpretation of quantum mechanics emphasizes the idea that this theory, whether it deals with pure or mixed states, does not govern individual systems but statistical ensembles of systems”. But it describes statistical ensembles only if the formal probabilities are taken to be observed frequentist probabilities, which requires Born’s rule. Thus the derivation is circular. The problem is still present in their recent paper [4, p.254], since they assume the maximum entropy principle, which (as discussed above in Subsection 1.3) requires the condensed form (BR-C) of Born’s rule. Moreover, to go from an ensemble to an individual case, they need an additional assumption called Postulate P. This postulate, ”although intuitive, is not a consequence of the mere quantum principles and has a macroscopic nature. We will accept its inclusion in quantum theory as a postulate (motivated by the macroscopic size of M)”. In effect they show that (BR-C) assumed for quantum statistical mechanics, together with their new Postulate P, implies that the measurement device they analyzed responds as the Born rule requires. Thus they reduced the measurement problem to their Postulate P.
@@
The main problem behind these difficulties is the need to find a rigorous derivation of a dissipative classical dynamics for a single pointer (or switch) without a silent recourse to the Born rule. A customary way to proceed (following Möhring & Smilansky [51]) is to start with the influence functional approach of Feynman & Vernon [27]. However, this approach has the Born rule built in – namely into the derivation (p.122) of the basic functional integral representation of the transition probabilities.
A solution of the measurement problem is indeed impossible unless one has a precise mathematical definition of when and how a mathematical model of a single quantum detector produces a definite response. E.g., Busch & Lahti [15, p.375] write: ”There is no consistent quantum measurement theory, unless an appropriate reinterpretation of what it means for an observable to have a definite value can be found.” This is the reason why, 100 years after the inception of quantum mechanics in terms of states and observables, the quantum measurement problem is still unsolved. It is the missing piece in the foundations of quantum mechanics.